OpenStudy (anonymous):

When doing absolute value, do you always have to isolate whatever is in the bars on one side of the equation?

OpenStudy (anonymous):

For example, |x+1|-4=-1. Is it absolutely necessary for me to isolate |x+1| on one side of the equal sign to solve the equation?

myininaya (myininaya):

If you are solving for x, you may have some operations to undo to get to x. so if you have |x|=3 and we are solving for x first recall |x| is sqrt(x^2) sqrt(x^2)=3 square both sides x^2=9 not square root both sides x=pm 3

myininaya (myininaya):

you don't have to do that first but you are trying to isolate x

myininaya (myininaya):

so eventually you will have to do it anyways

OpenStudy (anonymous):

I don't understand. I tried that equation without isolating the bar part and I only got it right after I isolated that part. In other equations involving absolute value, I usually just remove the bars and work the equation and I get it right. With this one I was only getting one of the right answers.

myininaya (myininaya):

\[|x+1|=\pm (x+1)\]

myininaya (myininaya):

you have two equations to solve

myininaya (myininaya):

\[|x+1|-4=-1\] \[\pm(x+1)-4=-1\] \[x+1-4=-1 \text{ or } -(x+1)-4=-1\]

myininaya (myininaya):

\[x-3=-1 \text{ or } -x-1-4=-1\] \[x=-1+3 \text{ or } -x-5=-1\]

myininaya (myininaya):

\[x=2 \text{ or } -x=4\]

myininaya (myininaya):

\[x=2 \text{ or } x=-4\]

myininaya (myininaya):

Time for checking! x=2 |2+1|-4=|3|-4=3-4=-1 GOOD! x=-4 |-4+1|-4=|-3|-4=3-4=-1 GOOD!

myininaya (myininaya):

Both of our solutions checked out!

OpenStudy (anonymous):

:) thanks a lot